# Learning constitutive models and rheology from partial flow measurements

### v1

@@ Abstract
Constitutive laws are at the core of fluid mechanics, relating the fluid stress to its deformation rate. Unlike Newtonian fluids, most industrial and biological fluids are non-Newtonian, exhibiting a nonlinear relation. Accurately characterizing this nonlinearity is essential for predicting flow behavior in real-world engineering and translational applications. Yet current methods fall short by relying on bulk rheometer data and simple fits that fail to capture behaviors relevant in complex geometries and flow conditions. Data-driven approaches can capture more complex behaviors, but lack interpretability or consistency. To close this gap, we leverage automatic differentiation to build an end-to-end framework for robust rheological learning. We develop a differentiable non-Newtonian fluid solver with a tensor basis neural network closure that learns stress directly from arbitrary flow measurements, such as velocimetry data. In parallel, we implement differentiable versions of major constitutive relations, enabling Bayesian model parametrization and selection from rheometer data. Our framework predicts flows in unseen geometries and ensures physical consistency and interpretability by matching neural network responses to known constitutive laws. Ultimately, this work lays the groundwork for advanced digital rheometry capable of comprehensively characterizing non-Newtonian and viscoelastic fluids under realistic in-situ or in-line operating conditions.
@@
Modeling and understanding how fluids behave in different conditions and geometries requires an accurate mathematical description of the relationship between local stress and local deformation. Such constitutive laws serve as a pillar of fluid mechanics, material science, and engineering, with applications from drug delivery to oil extraction [1, 2]. In its most basic form, a constitutive law gives a one-dimensional functional relationship between a scalar stress and a scalar strain rate [3, 4, 5, 6]. In many instances, such simple descriptions can be extended to three dimensions to provide relationships between tensorial quantities. More complex constitutive laws can involve additional evolution equations for stress tensors [7, 8, 9]. In such cases, the constitutive equations must be solved concurrently with the Navier–Stokes equations to obtain predictive models of fluid behavior [10, 11, 12].
@@
However, such predictions can often fail, highlighting a fundamental disconnect between how constitutive models are parameterized and how they are applied. For example, material parameters for a Boger fluid fit from simple shear rheometry often fail to predict the fluid’s strong extensional response in contractional geometries, sometimes predicting the opposite sign for the pressure drop [17, 18, 19, 20]. The cause of such failures are twofold. First, the choice of a constitutive law is itself an assumption. Parameters fit in one flow regime, such as simple shear, may mask the model’s inability to represent the physics of another, like mixed shear and extension [21, 22]. Second, standard rheological measurements, such as frequency sweeps and small-amplitude oscillatory tests, may lack sufficient kinematic detail required to constrain parameters or to distinguish between competing constitutive models [13, 23, 24, 25]. Consequently, a model may appear well-fit to rheometer data yet fail in predicting responses in other systems, not because its parameters are wrong, but because the underlying model form or the data used to fit it were not representative of the target application.
These challenges have motivated a growing body of work that moves away from classical constitutive models toward data-driven approaches for representing complex fluid behavior. One class of these methods has focused on modeling bulk rheometer data using machine-learning, including Rheological Universal Differential Equations (RUDEs) [26] and Rheology Informed Neural Networks (RhINNs) [27, 28, 29, 30, 31]. In these works, the authors replace components of classical constitutive laws with neural networks trained on experimental data, which are then used to predict stresses in unseen flow conditions. Other techniques, based on sparse symbolic regression, aim to discover simple, interpretable models directly from the data [32, 33]. These data-driven methods have shown promise in fitting small data sets to produce accurate stress predictions under certain conditions. However, the machine-learning models may still struggle to generalize when only limited bulk rheometry data are available for training, especially given the large number of parameters that must be inferred from sparse, bulk measurements.
@@
We first apply this procedure to the TBNN trained on the constriction–expansion flow of Fig. 2. Using the digital rheometer (Methods Sec. .4, we impose sinusoidal shear-rate forcings with amplitude $f=10$ and frequency $\omega=1$ and record the corresponding stress response predicted by the TBNN. We then fit both Newtonian and Carreau–Yasuda (CY) models to the resulting stress curves (Fig. 4a). The Newtonian model fails to reproduce the nonlinear features of the response, whereas the CY model captures them accurately, yielding parameter values that closely match those used in the original simulation (first row of Table 1). Comparison of the Bayesian Information Criterion (BIC) further confirms that the CY model provides a substantially better statistical description of the learned constitutive behavior than the Newtonian model, despite having more parameters.
@@
Run Source Carreau–Yasuda parameters $\eta_{s}$ $\Delta\mathrm{BIC}$
@@
While this may seem to be cause for concern, it indicates an important factor in Bayesian analysis, which is that our predictions are conditioned on the particular choice of forcing function. In this case, we generally followed the protocol for fitting rheometer-like data laid out in other data-driven approaches [26] while adding additional amplitudes to extend the range of shear stresses observed in the data. For this particular set of forcings, that the underlying constitutive behavior is indistinguishable from other models, illustrating the importance of experimental design and how insufficiently informative forcings can make distinct models indistinguishable.
@@
For any non-linear fitting approach, choosing the correct hyperparameters is critically important to achieve good convergence. When we repeated the Carrea-Yasuda parameter identification experiment on a new set of samples with slower hyperparameter (as described in Methods) the median error factor decreased to at most 1.003 for all parameters except for $\eta$. Our differentiable approach still systematically overestimated $\eta$ due to the lack of high shear rates sampled in the data set.
@@
$\overset{\kern 2.5pt\smalltriangledown}{\bm{\mathsf{\tau}}}+\lambda\bm{\mathsf{\tau}}=2\eta_{p}\bm{\mathsf{D}},$ (7)
@@
$\overset{\kern 2.5pt\smalltriangledown}{\bm{\mathsf{\tau}}}=\frac{\partial\bm{\mathsf{\tau}}}{\partial t}+\mathbf{u}\bm{\cdot}\bm{\nabla}\bm{\mathsf{\tau}}-\bm{\mathsf{\tau}}\bm{\cdot}\bm{\nabla}\mathbf{u}-\bm{\nabla}\mathbf{u}^{\mathrm{T}}\bm{\cdot}\bm{\mathsf{\tau}},$ (8)
@@
$\overset{\kern 2.5pt\smalltriangledown}{\bm{\mathsf{\tau}}}+\lambda\bm{\mathsf{\tau}}+\frac{\alpha\lambda}{\eta_{p}}\bm{\tau}\bm{\cdot}\bm{\tau}=2\eta_{p}\bm{\mathsf{D}},$ (9)
@@
$\overset{\kern 2.5pt\smallsquare}{\bm{\mathsf{\tau}}}+\left(1+\frac{\varepsilon\lambda}{\eta_{p}}\mathrm{tr}(\bm{\mathsf{\tau}})\right)\bm{\mathsf{\tau}}=2\eta_{p}\bm{\mathsf{D}}.$ (10)
@@
$\overset{\kern 2.5pt\smallsquare}{\bm{\mathsf{\tau}}}=\overset{\kern 2.5pt\smalltriangledown}{\bm{\mathsf{\tau}}}+\zeta(\bm{\mathsf{\tau}}\bm{\cdot}\bm{\mathsf{D}}+\bm{\mathsf{D}}\bm{\cdot}\bm{\mathsf{\tau}})$ (11)
@@
To mimic experimental measurements in Figs. S4S6, we apply a PIV observation operator that first performs Hann-windowed, separable smoothing over interrogation windows and then samples at vector-center locations:
@@
Given that the shear rheometer forms the backbone of most laboratory experiments in complex fluids, we first explore how differentiable solvers enable fitting models to bulk rheometer data, in line with previous data-driven approaches to the same topic [26]. In particular, we assume that the Weissenberg number is small enough that the flow in the rheometer gap becomes approximately two dimensional and spatially homogeneous. The result is that constitutive laws can be reduced to ordinary differential equations only in terms of time that are expressed terms of the rate of strain tensor $\bm{\mathsf{D}}$ and the stress $\bm{\mathsf{\sigma}}$, with no direct appearance of the flow $\mathbf{u}$.
Under the quasi-two-dimensional flow assumption of a shear rheometer, the gradient of the flow $\bm{\nabla}\mathbf{u}=\dot{\gamma}(t)\mathbf{e}_{12}$ only has one non-zero component in the shear direction characterized by a potentially time-dependent shear rate $\dot{\gamma}$, which is spatially homogeneous across the entire rheometer. In a shear rheometer, this shear rate $\dot{\gamma}$ may be prescribed and the resulting shear stress $\sigma_{12}$ measured by averaging the torsional resistance encountered by the top plate. Alternatively, the shear stress may be prescribed and the strain measured directly. We will use the former case, though the approach would work equally well in the latter case. While the shear strain is assumed to be unidirectional, the stress $\bm{\mathsf{\sigma}}$ has six independent components which can be non-zero due to non-linear coupling between the different stress terms in viscoelastic constitutive laws.
@@
where $\sigma_{12}$ is our predicted shear stress and $\theta$ the vector of material parameters that defines the model. Thus, fitting a given model to the data takes the form of a non-linear minimization problem over $\theta$, where the minimization must be performed through the solution to the ODE which governs $\sigma_{12}$ for a given $\dot{\gamma}$.
We perform the integration using a differentiable solver Diffrax [45], which allows us to automatically take gradients of the loss function, Eqn (34), with respect to the parameters $\theta$. We can use gradient-descent algorithms to efficiently find a set of parameters $\theta$ that best describe the data by minimizing the loss.
@@
where $\mathcal{L}$ is the mean squared error defined in Eqn (34). The Bayesian information criterion for a particular fit to a model is then
@@ Acknowledgments
We first examine the effect of spatial resolution by varying $W_{\mathrm{win}}$ while keeping the noise parameters fixed. As shown in Fig. S4a, the trained TBNN maintains its predictive accuracy, quantified by the relative root-mean-squared error (RRMSEu), even when the resolution is reduced to $13\times 13$ interrogation windows. This robustness demonstrates that spatially coarse flow measurements remain highly informative, containing sufficient kinematic diversity to recover the underlying constitutive behavior, far beyond what is accessible from traditional bulk rheometry. Example downsampled reconstructions are shown in Fig. S4b,c.
We next examine the effect of measurement noise using the same correlated, heteroskedastic model. Here the noise level on the horizontal axis of Fig. S5a is defined as the amplitude of the base scale $\sigma_{\mathrm{base}}=\alpha\,U_{95}$, expressed as a percentage of the 95th-percentile velocity magnitude $U_{95}$. As the noise amplitude increases, the RRMSE between the TBNN prediction and the ground-truth field rises slightly but remains low, even for perturbations as large as 4%  $U_{95}$. Although the total loss cannot decrease as much in these cases, reflecting the irreducible mismatch introduced by measurement noise (Fig. S6a), the trained model still captures the underlying flow structure with high fidelity (Fig. S6b). This robustness underscores a key advantage of our differentiable formulation: by enforcing the governing equations as hard constraints, the model resists overfitting noisy data and instead converges to the physically consistent constitutive relation, even when the available information is degraded.

### v2

@@ Abstract
Constitutive laws are at the core of fluid mechanics, relating the fluid stress to its deformation rate. Unlike Newtonian fluids, most industrial and biological fluids are non-Newtonian, exhibiting a nonlinear relation. Accurately characterizing this nonlinearity is essential for predicting flow behavior in real-world engineering and translational applications. Yet current methods fall short by relying on bulk rheometer data and simple fits that fail to capture behaviors relevant in complex geometries and flow conditions. Data-driven approaches can capture more complex behaviors, but lack interpretability or consistency. To close this gap, we leverage automatic differentiation to build an end-to-end framework for robust rheological learning. We develop a differentiable non-Newtonian fluid solver with a frame-invariant tensor basis neural network closure that learns stress directly from arbitrary flow measurements, such as velocimetry data. In parallel, we implement differentiable versions of major constitutive relations, enabling Bayesian model parametrization and selection from rheometer data. Our framework predicts flows in unseen geometries and ensures physical consistency and interpretability by matching neural network responses to known constitutive laws. Ultimately, this work lays the groundwork for advanced digital rheometry capable of comprehensively characterizing non-Newtonian and viscoelastic fluids under realistic in-situ or in-line operating conditions.
@@
Modeling and understanding how fluids behave in different conditions and geometries requires an accurate mathematical description of the relationship between local stress and local deformation. Such constitutive laws serve as a pillar of fluid mechanics, materials science, and engineering, with applications ranging from drug delivery to oil extraction [24, 34, 47, 14, 49, 15]. In its most basic form, a constitutive law gives a one-dimensional functional relationship between a scalar stress and a scalar strain rate [3, 6, 48, 18]. In many cases, such simple descriptions can be extended to three dimensions to provide relationships between tensorial quantities. More complex constitutive laws can involve additional evolution equations for stress tensors [35, 36, 16]. In such cases, the constitutive equations must be solved concurrently with the Navier–Stokes equations to obtain predictive models of fluid behavior [32, 28, 10].
@@
However, such predictions can often fail, highlighting a fundamental disconnect between how constitutive models are parameterized and how they are applied. For example, material parameters for a Boger fluid fit from simple shear rheometry often fail to predict the fluid’s strong extensional response in contractional geometries, sometimes predicting the opposite sign for the pressure drop [40, 20, 5, 19]. The causes of such failures are twofold. First, the choice of a constitutive law is itself an assumption. Parameters fit in one flow regime, such as simple shear, may mask the model’s inability to represent the physics of another, like mixed shear and extension [9, 12]. Second, standard rheological measurements, such as frequency sweeps and small-amplitude oscillatory tests, may lack sufficient kinematic detail required to constrain parameters or to distinguish between competing constitutive models [11, 33, 37, 17]. Consequently, a model may appear well-fit to rheometer data yet fail in predicting responses in other systems, not because its parameters are wrong, but because the underlying model form or the data used to fit it were not representative of the target application.
These challenges have motivated a growing body of work that moves away from classical constitutive models toward data-driven approaches for representing complex fluid behavior. One class of these methods has focused on modeling bulk rheometer data using machine learning, including Rheological Universal Differential Equations (RUDEs) [25] and Rheology Informed Neural Networks (RhINNs) [41, 8, 1, 30, 31]. In these works, the authors replace components of classical constitutive laws with neural networks trained on experimental data, which are then used to predict stresses in unseen flow conditions. Other techniques, based on sparse symbolic regression, aim to discover simple, interpretable models directly from the data [43, 42]. These data-driven methods have shown promise in fitting small data sets to produce accurate stress predictions under certain conditions. However, the machine learning models may still struggle to generalize when only limited bulk rheometry data are available for training, especially given the large number of parameters that must be inferred from sparse, bulk measurements.
@@
We first apply this procedure to the TBNN trained on the constriction–expansion flow of Fig. 2. Using the digital rheometer (Methods Sec. .4), we impose sinusoidal shear-rate forcings with amplitude $f=10$ and frequency $\omega=1$ and record the corresponding stress response predicted by the TBNN. We then fit both Newtonian and Carreau–Yasuda (CY) models to the resulting stress curves (Fig. 4a). The Newtonian model fails to reproduce the nonlinear features of the response, whereas the CY model captures them accurately, yielding parameter values that closely match those used in the original simulation (first row of Table 1). Comparison of the Bayesian Information Criterion (BIC) further confirms that the CY model provides a substantially better statistical description of the learned constitutive behavior than the Newtonian model, despite having more parameters.
@@
Run Source Carreau–Yasuda parameters $\eta_{N}$ $\Delta\mathrm{BIC}$
@@
While this may seem to be cause for concern, it indicates an important factor in Bayesian analysis, which is that our predictions are conditioned on the particular choice of forcing function. In this case, we generally followed the protocol for fitting rheometer-like data laid out in other data-driven approaches [25] while adding additional amplitudes to extend the range of shear stresses observed in the data. For this particular set of forcings, the underlying constitutive behavior is indistinguishable from other models, illustrating the importance of experimental design and how insufficiently informative forcings can make distinct models indistinguishable.
@@
For any nonlinear fitting approach, choosing the correct hyperparameters is critically important to achieve good convergence. When we repeated the Carreau-Yasuda parameter identification experiment on a new set of samples with slower hyperparameter (as described in Methods) the median error factor decreased to at most 1.003 for all parameters except for $\eta$. Our differentiable approach still systematically overestimated $\eta$ due to the lack of high shear rates sampled in the data set.
@@
$\overset{\kern 2.04861pt\smalltriangledown}{\bm{\mathsf{\tau}}}+\lambda\bm{\mathsf{\tau}}=2\eta_{p}\bm{\mathsf{D}},$ (7)
@@
$\overset{\kern 2.04861pt\smalltriangledown}{\bm{\mathsf{\tau}}}=\frac{\partial\bm{\mathsf{\tau}}}{\partial t}+\mathbf{u}\bm{\cdot}\bm{\nabla}\bm{\mathsf{\tau}}-\bm{\mathsf{\tau}}\bm{\cdot}\bm{\nabla}\mathbf{u}-\bm{\nabla}\mathbf{u}^{\mathrm{T}}\bm{\cdot}\bm{\mathsf{\tau}},$ (8)
@@
$\overset{\kern 2.04861pt\smalltriangledown}{\bm{\mathsf{\tau}}}+\lambda\bm{\mathsf{\tau}}+\frac{\alpha\lambda}{\eta_{p}}\bm{\tau}\bm{\cdot}\bm{\tau}=2\eta_{p}\bm{\mathsf{D}},$ (9)
@@
$\overset{\kern 2.04861pt\smallsquare}{\bm{\mathsf{\tau}}}+\left(1+\frac{\varepsilon\lambda}{\eta_{p}}\mathrm{tr}(\bm{\mathsf{\tau}})\right)\bm{\mathsf{\tau}}=2\eta_{p}\bm{\mathsf{D}}.$ (10)
@@
$\overset{\kern 2.04861pt\smallsquare}{\bm{\mathsf{\tau}}}=\overset{\kern 2.04861pt\smalltriangledown}{\bm{\mathsf{\tau}}}+\zeta(\bm{\mathsf{\tau}}\bm{\cdot}\bm{\mathsf{D}}+\bm{\mathsf{D}}\bm{\cdot}\bm{\mathsf{\tau}})$ (11)
@@
To mimic experimental measurements, in Figs. S4S6, we apply a PIV observation operator that first performs Hann-windowed, separable smoothing over interrogation windows and then samples at vector-center locations:
@@
Given that the shear rheometer forms the backbone of most laboratory experiments in complex fluids, we first explore how differentiable solvers enable fitting models to bulk rheometer data, in line with previous data-driven approaches to the same topic [25]. In particular, we assume that the Weissenberg number is small enough that the flow in the rheometer gap becomes approximately two-dimensional and spatially homogeneous. The result is that constitutive laws can be reduced to ordinary differential equations only in terms of time that are expressed in terms of the rate of strain tensor $\bm{\mathsf{D}}$ and the stress $\bm{\mathsf{\sigma}}$, with no direct appearance of the flow $\mathbf{u}$.
Under the quasi-two-dimensional flow assumption of a shear rheometer, the gradient of the flow $\bm{\nabla}\mathbf{u}=\dot{\gamma}(t)\mathbf{e}_{12}$ only has one non-zero component in the shear direction characterized by a potentially time-dependent shear rate $\dot{\gamma}$, which is spatially homogeneous across the entire rheometer. In a shear rheometer, this shear rate $\dot{\gamma}$ may be prescribed and the resulting shear stress $\sigma_{12}$ measured by averaging the torsional resistance encountered by the top plate. Alternatively, the shear stress may be prescribed and the strain measured directly. We will use the former case, though the approach would work equally well in the latter case. While the shear strain is assumed to be unidirectional, the stress $\bm{\mathsf{\sigma}}$ has six independent components which can be non-zero due to nonlinear coupling between the different stress terms in viscoelastic constitutive laws.
@@
where $\sigma_{12}$ is our predicted shear stress and $\theta$ the vector of material parameters that defines the model. Thus, fitting a given model to the data takes the form of a nonlinear minimization problem over $\theta$, where the minimization must be performed through the solution to the ODE which governs $\sigma_{12}$ for a given $\dot{\gamma}$.
We perform the integration using a differentiable solver Diffrax [21], which allows us to automatically take gradients of the loss function, Eq. (34), with respect to the parameters $\theta$. We can use gradient-descent algorithms to efficiently find a set of parameters $\theta$ that best describe the data by minimizing the loss.
@@
where $\mathcal{L}$ is the mean squared error defined in Eq. (34). The Bayesian information criterion for a particular fit to a model is then
@@ Acknowledgments
We first examine the effect of spatial resolution by varying $W_{\mathrm{win}}$ while keeping the noise parameters fixed. As shown in Fig. S4a, the trained TBNN maintains its predictive accuracy, quantified by the relative root-mean-squared error (RRMSEu), even when the resolution is reduced to $13\times 13$ interrogation windows. This robustness demonstrates that spatially coarse flow measurements remain highly informative, containing sufficient kinematic diversity to recover the underlying constitutive behavior far beyond what is accessible from traditional bulk rheometry. Example downsampled reconstructions are shown in Fig. S4b,c.
We next examine the effect of measurement noise using the same correlated, heteroskedastic model. Here the noise level on the horizontal axis of Fig. S5a is defined as the amplitude of the base scale $\sigma_{\mathrm{base}}=\alpha\,U_{95}$, expressed as a percentage of the 95th-percentile velocity magnitude $U_{95}$ (computed on the $29\times 29$ downsampled vector grid, where the noise is added). As the noise amplitude increases, the RRMSE between the TBNN prediction and the ground-truth field rises slightly but remains low, even for perturbations as large as 4%  $U_{95}$. Although the total loss cannot decrease as much in these cases, reflecting the irreducible mismatch introduced by measurement noise (Fig. S6a), the trained model still captures the underlying flow structure with high fidelity (Fig. S6b). This robustness underscores a key advantage of our differentiable formulation: by enforcing the governing equations as hard constraints, the model resists overfitting noisy data and instead converges to the physically consistent constitutive relation, even when the available information is degraded.